# Complexities of Theories of Profinite Subgroups of $S_{\omega}$ via Tree Presentations

### v1

@@ Abstract
Although $S_{\omega}$ (the group of all permutations of $\mathbb{N}$) is size continuum, both it and its closed subgroups can be presented as the set of paths through a countable tree. The subgroups of $S_{\omega}$ that can be presented this way with finite branching trees are exactly the profinite ones. We use these tree presentations to find upper bounds on the complexity of the existential theories of profinite subgroups of $S_{\omega}$, as well as to prove sharpness for these bounds. These complexity results enable us to distinguish a simple subclass of profinite groups, those with orbit independence, for which we find an upper bound on the complexity of the entire first order theory.
@@ 1 Introduction
Traditional computable structure theory deals only with countable structures. As a result, it cannot be used to study most subgroups of $S_{\omega}$ (the group of all permutations of $\mathbb{N}$). However, as described in Section 2, a large class of these subgroups (specifically the closed subgroups) can be presented as the set of paths through a countable tree. We will focus our attention on the subgroups that can be presented as the paths through a finite branching tree (the compact subgroups), which are exactly the profinite ones. In [5] the author uses such a presentation to study the absolute Galois group of $\mathbb{Q}$, which is indeed a profinite group that can be viewed as a subgroup of $S_{\omega}$ after fixing an enumeration of the algebraic closure of $\mathbb{Q}$.
@@
Much interest in profinite groups stems from their connection to Galois theory. As shown in [8], every profinite group is the Galois group of some field extension. However, the purpose of this paper is to examine profinite groups simply as groups in their own right. It can be difficult to get an effective handle on uncountable groups, but when such a group acts on $\mathbb{N}$ by permutations we are given the opportunity to do so. Thus, we restrict our attention to the profinite subgroups of $S_{\omega}$.
@@
Effective notions for profinite groups within the context of Galois theory were examined in [4] and further in [3]. Following this work, effective notions for profinite groups in general were studied in [7]. The authors of [3] and [7] define a profinite group to be recursively profinite if it is isomorphic to the inverse limit of a uniformly computable sequence of finite groups and surjective homomorphisms. As we will see in Proposition 4, a profinite group $P$ is recursively profinite if and only if it is isomorphic to a subgroup $G$ of $S_{\omega}$ such that $T_{G}$ (the tree that represents $G$) is computable. More recently, effectively closed subgroups of $S_{\omega}$ were examined in [2].
@@
In Section 3, we use tree presentations to determine bounds on the complexity of the existential theory of profinite subgroups of $S_{\omega}$. Note that since all elements of $S_{\omega}$ are functions from $\mathbb{N}\to\mathbb{N}$, an existential sentence about a subgroup is a $\Sigma^{1}_{1}$ statement. However, we will see that the existential theory of any profinite subgroup $G$ of $S_{\omega}$ is $\Sigma^{0}_{2}$ relative to the degree of $T_{G}$. Additionally, if $G$ has orbit independence, then the existential theory is $\Sigma^{0}_{1}$ relative to the degree of $T_{G}$. We will also show that these bounds are sharp. Specifically, there exists a profinite $G$ with orbit independence such that $T_{G}$ is computable and the existential theory of $G$ is $\Sigma^{0}_{1}$ complete, and such a $G$ without orbit independence such that the existential theory is $\Sigma^{0}_{2}$ complete. Last, we show that the (entire) first order theory of a profinite $G$ with orbit independence is $\Delta^{0}_{2}$ relative to the degree of $T_{G}$.
@@ 2 Tree Presentations
Let $G$ be a subgroup of $S_{\omega}$. We define the tree $T_{G}$ to be the subtree of $\mathbb{N}^{<\omega}$ containing all initial segments of elements of $G$. That is,
$T_{G}:=\{\tau\in\mathbb{N}^{<\omega}:(\exists g\in G,n\in\mathbb{N})[\tau=g(0)g(1)\cdots g(n)]\}$
@@
Let $G$ be a subgroup of $S_{\omega}$. We define the degree of $T_{G}$ to be the join of the Turing degrees of
@@
$i(g)=g(0)g(1)\cdots$
[…]
@@
Let $G$ be a subgroup of $S_{\omega}$. The map $i:G\to[T_{G}]$ is a bijection if and only if $i(G)$ is a closed subset of $[T_{G}]$. ∎
We say that $G$ is a closed group when $i(G)$ is closed. Thus, the subgroups of $S_{\omega}$ that can be represented as the paths through this type of tree are exactly the closed ones. Additionally, we say that $G$ is a compact group if $i(G)$ is compact.
@@
Given a subgroup $G$ of $S_{\omega}$ and $n\in\mathbb{N}$, we define the orbit of $n$ under $G$ as
@@
Let $G$ be a subgroup of $S_{\omega}$. The following are equivalent:
@@
Suppose that $G$ is compact . Since our topology is Hausdorff, we have that $G$ is closed. By Proposition 1, we have that $i(G)=[T_{G}]$. Assume towards a contradiction that there is some $n\in\mathbb{N}$ with $\text{orb}_{G}(n)$ infinite. Let $\{\tau_{i}\}_{i\in\mathbb{N}}$ be the (infinite) collection of all elements of $T_{G}$ of length $n+1$. Note that $\{\{f:\tau_{i}\sqsubset f\}\}_{i\in\mathbb{N}}$ is an open cover of $[T_{G}]=i(G)$ with no finite subcover, which contradicts that $G$ is compact. Hence, $(1)\implies(2)$.
@@
We have that all profinite subgroups of $S_{\omega}$ will have countably many orbits (all of which are finite). We fix an enumeration of these orbits as follows:
@@
Let $G$ be a profinite subgroup of $S_{\omega}$. Define $\{O_{G,i}\}_{i\in\mathbb{N}}$ so that $O_{G,0}=\text{orb}_{G}(0)$ and $O_{G,n+1}$ is the orbit of the least natural number not in any $O_{G,m}$ with $m\leq n$.
@@
Let $G$ be a profinite subgroup of $S_{\omega}$. Given $g\in G$ and $k\in\mathbb{N}$, define $g_{k}=g\upharpoonright\bigcup_{i\leq k}O_{G,i}$. Define
@@
Let $G$ be a profinite subgroup of $S_{\omega}$. Define
@@
Let $G$ be a profinite subgroup of $S_{\omega}$. We say that $G$ has orbit independence if it is isomorphic to the Cartesian product of all $H_{k}$. That is,
@@
For an example of a profinite subgroup that does not have orbit independence, consider
[…]
where $1_{G}$ denotes the identity permutation. Note that $H_{0}\cong H_{1}\cong C_{2}$ (the cyclic group on $2$ elements) and $H_{n}$ is trivial for all $n>1$. Thus, $\prod H_{k}\cong C_{2}\times C_{2}$ but $G\cong C_{2}$. As we shall see, groups with orbit independence tend to be simpler to work with.
It should be noted that our definition for the degree of $T_{G}$ is compatible with the notion of a recursively profinite group used by La Roche and Smith.
@@
A profinite group $P$ is recursively profinite if and only if it is isomorphic to a subgroup $G$ of $S_{\omega}$ with $T_{G}$ computable.
@@
For the other direction, suppose that $\{P_{n},\pi_{n}\}_{n\in\mathbb{N}}$ is as in definition 10. For each $n\in\mathbb{N}$, let $N_{n}$ be a natural number such that $P_{n}$ is isomorphic to a subgroup of $S_{N_{n}}$ (the group of permutations of $\{0,...,N_{n}-1\}$). Define $f_{0}:P_{0}\to S_{N_{0}}$ such that $f_{0}$ is a group embedding. Given $f_{n}$, define $f_{n+1}:P_{n+1}\to S_{N_{n+1}}$ such that $f_{n+1}$ is a group embedding and “respects” $\pi_{n+1}$ in the sense that if $\pi_{n+1}(p_{n+1})=p_{n}$, then $f_{n+1}(p_{n+1})\upharpoonright N_{n}=f_{n}(p_{n})$. Define $G$ to be the set of $g\in S_{\omega}$ such that for all $n\in\mathbb{N}$, there exists a $p\in P_{n}$ with $f_{n}(p)=g\upharpoonright N_{n}$. We have that $G\cong P$ and that $T_{G}$ is exactly the set
@@
It is clear that the domain and branching function of $T_{G}$ are computable, hence we have that $T_{G}$ is computable. ∎
@@ 3 Complexity of Theories
We now consider the complexity of the existential theory of a profinite subgroup of $S_{\omega}$. To do so, we must first establish a few lemmas.
@@
A positive formula is a first order formula that can be expressed without the use of any negation symbols. A negative formula is the negation of a positive formula.
@@
Let $G$ be a profinite subgroup of $S_{\omega}$ and let $\alpha^{+}$ be a positive formula in the language of groups. If $k<l$, then
@@
If $\alpha^{+}$ is quantifier free, than it is expressible as a disjunction of conjunctions of atomic formulas. Recall that an atomic formula in the language of groups is equivalent to the statement that some word is equal to the identity. If $G_{l}\models\alpha^{+}(\bar{g}_{l})$, then one of the disjuncts must hold which just means that some collection of words $W_{1}(\bar{g}_{l}),...,W_{n}(\bar{g}_{l})$ over the alphabet $\{x,x^{-1}:x\in\bar{g}_{l}\}$ are all equal to $1_{G_{l}}$. Since each element of $\bar{g}_{k}$ is an initial segment of an element of $\bar{g}_{l}$, we must also have that $W_{1}(\bar{g}_{k})=\cdots=W_{n}(\bar{g}_{k})=1_{G_{k}}$ and so $G_{k}\models\alpha^{+}(\bar{g}_{k})$.
Now suppose that the result holds for all $\Sigma_{n}$ and $\Pi_{n}$ formulas. If $\alpha^{+}$ is $\Sigma_{n+1}$, then we have $\alpha^{+}\equiv\exists\bar{x}\beta^{+}$ where $\beta^{+}$ is a $\Pi_{n}$ positive formula. If $G_{l}\models(\exists\bar{x})\beta^{+}(\bar{g}_{l},\bar{x})$ then there exists some $\bar{h}\in G^{<\omega}$ such that $G_{l}\models\beta^{+}(\bar{g}_{l},\bar{h}_{l}).$ Since the result holds of $\Pi_{n}$ formulas, we get that $G_{k}\models\beta^{+}(\bar{g}_{k},\bar{h}_{k})$ and so $G_{k}\models\alpha^{+}(\bar{g}_{k})$.
If $\alpha^{+}$ is $\Pi_{n+1}$, then we have $\alpha^{+}\equiv\forall\bar{x}\beta^{+}$ where $\beta^{+}$ is $\Sigma_{n}$. If $G_{l}\models(\forall\bar{x})\beta^{+}(\bar{g}_{l},\bar{x})$ then $G_{l}\models\beta^{+}(\bar{g}_{l},\bar{h}_{l})$ for all $\bar{h}\in G^{<\omega}$. Again, since the result holds for all $\Sigma_{n}$ formulas we have that $G_{k}\models\beta^{+}(\bar{g}_{k},\bar{h}_{k})$ and so $G_{k}\models\alpha^{+}(\bar{g}_{k})$. ∎
@@
Let $G$ be a profinite subgroup of $S_{\omega}$. If $\alpha$ is quantifier free, then $G\models\alpha(\bar{g})$ if and only if $G_{k}\models\alpha(\bar{g}_{k})$ for all sufficiently large $k\in\mathbb{N}$.
@@
Let $\alpha\equiv\neg\beta$ with $\beta$ atomic. Suppose $G\models\neg\beta(\bar{g})$. This gives that there is a $k$ such that $G_{k}\models\neg\beta(\bar{g}_{k})$. By the previous lemma, we must have that $G_{l}\models\neg\beta(\bar{g}_{l})$ for all $l\geq k$ and thus for all sufficiently large $l$.
@@
Every atomic sentence in the language of groups is true in every group.
Proof
Every such sentence has the form
$1^{n}=1^{m}$
with $n,m\in\mathbb{Z}$.∎
The following lemma gives an example of the power of orbit independence.
Lemma 4
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. Let $\alpha$ be an existential sentence in the language of groups. We have that $G\models\alpha$ if and only if $G_{k}\models\alpha$ for some $k\in\mathbb{N}$.
@@
Now suppose that $G_{k}\models\alpha$. We have that for some $\bar{\gamma}\in G_{k}^{<\omega}$, $G_{k}\models\beta(\bar{\gamma})$. Define an embedding $f:G_{k}\to G$ such that $f(\gamma)\upharpoonright\text{dom}(\gamma)=\gamma$ and $f(\gamma)$ is just the identity on all orbits $O_{G,l}$ with $l>k$. Note that since $G$ has orbit independence, we will in fact have that $f(\gamma)\in G$.
Note that
$\beta(\bar{x})\equiv\bigvee_{i}\bigwedge_{j}\beta^{+}_{i,j}(\bar{x})\,\&\,\beta^{-}_{i,j}(\bar{x})$
where each $\beta^{+}$ is atomic and each $\beta^{-}$ negated atomic. Since $G_{k}\models\beta(\bar{\gamma})$, there is some $i$ such that $G_{k}\models\bigwedge_{j}\beta^{+}_{i,j}(\bar{\gamma})\,\&\,\beta^{-}_{i,j}(\bar{\gamma}).$ It is clear that any negated atomic formula that holds of $\bar{\gamma}$ in $G_{k}$ will also hold of $f(\bar{\gamma})$ in $G$. Thus if $G\nvDash\alpha$, there would have to be some $j$ such that $G\nvDash\beta^{+}_{i,j}(f(\bar{\gamma}))$. However since we are assuming $G_{k}\models\beta^{+}_{i,j}(\bar{\gamma})$, there would have to be some $l>k$ such that $H_{l}\models\neg\beta^{+}_{i,j}(f(\bar{\gamma})\upharpoonright O_{G,l})$. Since $\bar{\gamma}$ is just the identity on all orbits $O_{G,l}$ with $l>k$, we would have $H_{l}\models\neg\beta^{+}_{i,j}(\bar{1})$. The formula $\beta^{+}_{i,j}(\bar{1})$ is an atomic sentence in the language of groups, and so by Lemma 3 we get that $H_{l}\models\beta^{+}_{i,j}(\bar{1})$. Hence, we must have that $G\models\alpha$. ∎
@@
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. The existential theory of $G$ is $\Sigma^{0}_{1}$ relative to $\deg(T_{G})$.
@@
There exists a profinite subgroup $G$ of $S_{\omega}$ with orbit independence such that $T_{G}$ is computable and the existential theory of $G$ is $\Sigma^{0}_{1}$ complete.
@@
Let $G$ be any profinite subgroup of $S_{\omega}$ (not necessarily with orbit independence). The existential theory of $G$ is $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$.
@@
There exists a profinite subgroup $G$ of $S_{\omega}$ (without orbit independence) with $T_{G}$ computable such that the existential theory of $G$ is $\Sigma^{0}_{2}$ complete.
@@
Recall that the set $\text{Fin}=\{e\in\mathbb{N}:|W_{e}|<\infty\}$ (where $W_{e}$ is the domain of $\Phi_{e}$, the $e$th Turing program) is $\Sigma^{0}_{2}$ complete. Let $\{p_{n}\}_{n\in\mathbb{N}}$ be the sequence of all primes. Given $n\in\mathbb{N}$, define the formula
@@
This, along with Lemma 3.2, ensures that the existential theory of $G$ is $\Sigma^{0}_{2}$ complete.
@@
Let $G$ be a profinite subgroup of $S_{\omega}$ with orbit independence. Let $\alpha$ be any first order sentence in the language of groups. We have $G\models\alpha$ if and only if $G_{k}\models\alpha$ for all sufficiently large $k\in\mathbb{N}$.

### v2

# Elementarity of Subgroups and Complexity of Theories for Profinite Groups

@@ Abstract
Although $S_{\infty}$ (the group of all permutations of $\mathbb{N}$) is size continuum, both it and its closed subgroups can be presented as the set of paths through a countable tree. The subgroups of $S_{\infty}$ that can be presented this way with finite branching trees are exactly the profinite ones. We use these tree presentations to find upper bounds on the complexity of the existential theories of profinite subgroups of $S_{\infty}$, as well as to prove sharpness for these bounds. These complexity results enable us to distinguish a simple subclass of profinite groups, those with orbit independence, for which we find an upper bound on the complexity of the entire first order theory. Additionally, given a profinite subgroup $G$ of $S_{\infty}$ and a Turing ideal $I$ we define $G_{I}$ to be the set of elements in $G$ whose Turing degree lies in $I$. We examine to what extent and under what conditions $G_{I}$ will be an elementary subgroup of $G$. In particular, we construct a profinite group whose subgroup of computable elements is not elementary even for existential formulas.
@@ 1 Introduction
Traditional computable structure theory deals only with countable structures. As a result, it cannot be used to study most subgroups of $S_{\infty}$ (the group of all permutations of $\mathbb{N}$). However as described in Section 2, a large class of these subgroups (specifically the closed subgroups) can be presented as the set of paths through a countable tree, similarly to how the real numbers are presented in computable analysis [17]. We will focus our attention on the subgroups of $S_{\infty}$ that can be presented as the paths through a finite branching tree (the compact subgroups), which are exactly the profinite ones. In [12] Miller uses such a presentation to study the absolute Galois group of $\mathbb{Q}$, which is indeed a profinite group that can be viewed as a subgroup of $S_{\infty}$ after fixing an enumeration of the algebraic closure of $\mathbb{Q}$. A well known result from computable analysis is that the field of computable real numbers is a real closed field and thus an elementary substructure of the field of real numbers. In contrast, Theorem 4.1 gives a profinite group whose subgroup of computable elements is not an elementary subgroup.
@@
Much interest in profinite groups stems from their connection to Galois theory. All Galois groups are profinite, and every profinite group is the Galois group of some field extension [16]. However, the purpose of this paper is to examine profinite groups simply as groups in their own right. It can be difficult to get an effective handle on uncountable groups, but when such a group acts on $\mathbb{N}$ by permutations we are given the opportunity to do so. Thus, we restrict our attention to the profinite subgroups of $S_{\infty}$.
@@
Effective notions for profinite groups within the context of Galois theory were examined by Metakides and Nerode in [9] and further by La Roche in [7]. Following this work, effective notions for profinite groups in general were studied by Smith in [14]. The authors of [7] and [14] define a profinite group to be recursively profinite if it is isomorphic to the inverse limit of a uniformly computable sequence of finite groups and surjective homomorphisms. As we will see in Proposition 5, a profinite group $P$ is recursively profinite if and only if it is isomorphic to a subgroup $G$ of $S_{\infty}$ such that $T_{G}$ (the tree that represents $G$) is computable.
There have been many recent developments in the field of computable analysis and topology, much of which has been applied to profinite groups. In [2], Downey and Melnikov show that a profinite group is recursively profinite if and only if it can be presented as a computably compact Polish space in which the group operation is computable. This is further equivalent to having computable Haar measure, as shown by Pauly, Seon and Ziegler in [11]. As for abelian profinite groups, Melnikov shows in [8] that such a group is recursively profinite if and only if its Pontryagin dual has a computable presentation. In [5] Koh, Melnikov and Ng compare several definitions of the more general notion of computable topological group and profinite groups are used to illustrate some of the differences. In [4] Greenberg, Melnikov, Nies and Turetsky examine what they call “effectively closed” subgroups of $S_{\infty}$. This notion is not equivalent to being a closed subgroup $G$ of $S_{\infty}$ with $T_{G}$ computable, as the authors construct an effectively closed profinite subgroup of $S_{\infty}$ that is not recursively profinite. In particular, no tree presentation $T_{G}$ of the group can have computable branching.
@@
In Section 3 we use tree presentations to determine bounds on the complexity of the existential theory of profinite subgroups of $S_{\infty}$. Note that since all elements of $S_{\infty}$ are functions from $\mathbb{N}\to\mathbb{N}$, an existential sentence about a subgroup is a $\Sigma^{1}_{1}$ statement. However, we will see that the existential theory of any profinite subgroup $G$ of $S_{\infty}$ is $\Sigma^{0}_{2}$ relative to the degree of $T_{G}$. Additionally, if $G$ has orbit independence, then the existential theory is $\Sigma^{0}_{1}$ relative to the degree of $T_{G}$. We will also show that these bounds are sharp. Specifically, there exists a profinite $G$ with orbit independence such that $T_{G}$ is computable and the existential theory of $G$ is $\Sigma^{0}_{1}$ complete, and such a $G$ without orbit independence such that the existential theory is $\Sigma^{0}_{2}$ complete. Last, we show that the (entire) first order theory of a profinite $G$ with orbit independence is $\Delta^{0}_{2}$ relative to the degree of $T_{G}$.
In Section 4 we consider countable subgroups of profinite subgroups $G$ of $S_{\infty}$. In particular, given a Turing ideal $I$ we define $G_{I}$ to be the subgroup of $G$ whose elements (viewed as functions from $\mathbb{N}\to\mathbb{N}$) have Turing degree in $I$. We examine to what degree such a $G_{I}$ will be an elementary subgroup of $G$. As shown by Theorem 4.1, it possible to have that $G_{I}$ will not be elementary even for purely existential formulas. However, if either $G$ has orbit independence or $I$ is a Scott ideal, then $G_{I}$ will be elementary at least for existential formulas. Additionally if we have both that $G$ has orbit independence and that $I$ is a Scott ideal, then $G_{I}$ is an elementary subgroup of $G$.
@@ 2 Tree Presentations
Let $G$ be a subgroup of $S_{\infty}$. We define the tree $T_{G}$ to be the subtree of $\mathbb{N}^{<\omega}$ containing all initial segments of elements of $G$. That is,
$T_{G}:=\left\{\tau\in\mathbb{N}^{<\omega}:(\exists g\in G,n\in\mathbb{N})\left[\tau=g(0)g^{-1}(0)g(1)g^{-1}(1)\cdots g(n)g^{-1}(n)\right]\right\}$
@@
Let $G$ be a subgroup of $S_{\infty}$. We define the degree of $T_{G}$ to be the join of the Turing degrees of
@@
$i(g)=g(0)g^{-1}(0)g(1)g^{-1}(1)\cdots$
is an embedding. However, it is possible for there to be additional paths through $T_{G}$ that do not correspond to any element of $G$ in such a way. For example, consider the group $G$ generated by $\{(0\,1),(2\,3),(4\,5),$...\} where (nm)(n\,m) denotes the permutation that swaps $n$ and $m$ and leaves everything else fixed. We see that $G$ is countable but $T_{G}]$ is size continuum. The following proposition gives a simple topological condition for when a group $G$ corresponds nicely with $T_{G}]$.
@@
Let $G$ be a subgroup of $S_{\infty}$. The map $i:G\to[T_{G}]$ is a bijection if and only if $i(G)$ is a closed subset of $T_{G}]$. ∎
We say that $G$ is a closed group when $i(G)$ is closed. Thus, the subgroups of $S_{\infty}$ that can be represented as the paths through this type of tree are exactly the closed ones. Additionally, we say that $G$ is a compact group if $i(G)$ is compact.
@@
Given a subgroup $G$ of $S_{\infty}$ and $n\in\mathbb{N}$, we define the orbit of $n$ under $G$ as
@@
Let $G$ be a subgroup of $S_{\infty}$. The following are equivalent:
@@
Suppose that $G$ is compact. Since our topology is Hausdorff, we have that $G$ is closed. By Proposition 1, $i(G)=[T_{G}]$. Assume towards a contradiction that there is some $n\in\mathbb{N}$ with $\text{orb}_{G}(n)$ infinite. Let $\{\tau_{i}\}_{i\in\mathbb{N}}$ be the (infinite) collection of all elements of $T_{G}$ of length $n+1$. Note that $\{\{f:\tau_{i}\sqsubset f\}\}_{i\in\mathbb{N}}$ is an open cover of $T_{G}]=i(G)$ with no finite subcover, which contradicts that $G$ is compact. Hence, $(1)\implies(2)$.
@@
We have that all profinite subgroups of $S_{\infty}$ will have countably many orbits (all of which are finite). We fix an enumeration of these orbits as follows:
@@
Let $G$ be a profinite subgroup of $S_{\infty}$. Define $\{O_{G,i}\}_{i\in\mathbb{N}}$ so that $O_{G,0}=\text{orb}_{G}(0)$ and $O_{G,n+1}$ is the orbit of the least natural number not in any $O_{G,m}$ with $m\leq n$.
@@
Let $G$ be a profinite subgroup of $S_{\infty}$. Given $g\in G$ and $k\in\mathbb{N}$, define $g_{k}=g\upharpoonright\bigcup_{i\leq k}O_{G,i}$. Define
@@
Let $G$ be a profinite subgroup of $S_{\infty}$. Define
@@
Let $G$ be a profinite subgroup of $S_{\infty}$. We say that $G$ has orbit independence if it is isomorphic to the Cartesian product of all $H_{k}$. That is,
@@
It should be noted that the property of orbit independence is not preserved under isomorphism. In fact, every profinite group is isomorphic to a subgroup of $S_{\infty}$ without orbit independence. For an example, consider the groups
$G=\{1_{G},(0\,1)(2\,3)\}$
and
$G^{\prime}=\{1_{G^{\prime}},(0\,1)\}$
where $1_{G}$ denotes the identity permutation. Note that $G\cong G^{\prime}\cong\mathbb{Z}/2\mathbb{Z}$ and it is clear that $G^{\prime}$ has orbit independence. However for $G$, we have $H_{0}\cong H_{1}\cong\mathbb{Z}/2\mathbb{Z}$ and $H_{n}$ is trivial for all $n>1$. Thus, $\prod H_{k}\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}\ncong G$.
A profinite subgroup of $S_{\infty}$ is isomorphic to such a subgroup with orbit independence if and only if it is isomorphic to the direct product of finite groups. Since not all profinite groups can be expressed as such a direct product, some profinite groups have no orbit independent representation. This is illustrated by the following proposition:
Proposition 4
Let $G$ be a (nontrivial) profinite subgroup of $S_{\infty}$. If $G$ is torsion free, then $G$ does not have orbit independence.
Proof
Suppose we had $G\cong\prod_{k\in\mathbb{N}}H_{k}$. Take the least $n$ such that $H_{n}$ is nontrivial and let $h$ be some non-identity element of $H_{n}$. Since all $H_{k}$ are finite, $h$ is a torsion element. Define the element $(h_{k})_{k\in\mathbb{N}}\in G$ such that $h_{n}=h$ and $h_{m}$ is the identity when $m\neq n$. This would be a torsion element of $G$. ∎
Orbit independent profinite groups do arise naturally, including in the context of Galois theory. For example, consider the field $F=\mathbb{Q}(\sqrt{p_{i}})_{i\in\mathbb{N}}$ where $p_{i}$ denotes the $i$th prime number. We have that the absolute Galois group of $F$ is isomorphic to the direct product of countably many copies of the two element group, which is naturally embedded as an orbit independent subgroup of $S_{\infty}$.
Our definition for the degree of $T_{G}$ is compatible with the notion of a recursively profinite group used by La Roche and Smith.
@@
A profinite group $P$ is recursively profinite if and only if it is isomorphic to a subgroup $G$ of $S_{\infty}$ with $T_{G}$ computable.
@@
For the other direction, suppose that $\{P_{n},\pi_{n}\}_{n\in\mathbb{N}}$ is as in definition 10. For each $n\in\mathbb{N}$, let $N_{n}$ be a natural number such that $P_{n}$ is isomorphic to a subgroup of $S_{N_{n}}$ (the group of permutations of $\{0,...,N_{n}-1\}$). Define $f_{0}:P_{0}\to S_{N_{0}}$ such that $f_{0}$ is a group embedding. Given $f_{n}$, define $f_{n+1}:P_{n+1}\to S_{N_{n+1}}$ such that $f_{n+1}$ is a group embedding and “respects” $\pi_{n+1}$ in the sense that if $\pi_{n+1}(p_{n+1})=p_{n}$, then $f_{n+1}(p_{n+1})\upharpoonright N_{n}=f_{n}(p_{n})$. Define $G$ to be the set of $g\in S_{\infty}$ such that for all $n\in\mathbb{N}$, there exists a $p\in P_{n}$ with $f_{n}(p)=g\upharpoonright N_{n}$. We have that $G\cong P$ and that $T_{G}$ is exactly the set
@@
It is clear that the domain and branching function of $T_{G}$ are computable, hence $T_{G}$ is computable. ∎
@@ 3 Complexity of Theories
We now consider the complexity of the existential theory of a profinite subgroup of $S_{\infty}$. To do so, we must first establish a few lemmas.
@@
A positive formula is a first order formula that can be expressed using only the logical connectives $\wedge$ and $\vee$ without the use of negation symbols. A negative formula is the negation of a positive formula.
It should be noted that quantifiers are allowed in positive/negative formulas and that many formulas are neither positive nor negative. The following lemma illustrates a well known property of positive formulas.
@@
Positive formulas are preserved under surjective homomorphisms. That is, if $\alpha^{+}$ is a positive formula and $f:\mathcal{A}\to\mathcal{B}$ is a surjective homomorphism, then $\mathcal{A}\models\alpha^{+}(\bar{a})$ implies that $\mathcal{B}\models\alpha^{+}(f(\bar{a}))$. ∎
Lemma 2
Let $G$ be a profinite subgroup of $S_{\infty}$ and let $\alpha^{+}$ be a positive formula in the language of groups. If $k<l$, then
@@
The natural projection $\pi:G_{l}\to G_{k}$ defined such that $\pi(g_{l})=g_{k}$ is a surjective homomorphism. Hence, the result follows from Lemma 1. ∎
@@
Let $G$ be a profinite subgroup of $S_{\infty}$. If $\alpha$ is quantifier free, then $G\models\alpha(\bar{g})$ if and only if $G_{k}\models\alpha(\bar{g}_{k})$ for all sufficiently large $k\in\mathbb{N}$.
@@
Let $\alpha\equiv\neg\beta$ with $\beta$ atomic. Suppose $G\models\neg\beta(\bar{g})$. This gives that there is a $k$ such that $G_{k}\models\neg\beta(\bar{g}_{k})$. By the previous corollary, we must have that $G_{l}\models\neg\beta(\bar{g}_{l})$ for all $l\geq k$ and thus for all sufficiently large $l$.
@@
Let $G$ be a profinite subgroup of $S_{\infty}$ with orbit independence. Let $\alpha$ be an existential sentence in the language of groups. We have that $G\models\alpha$ if and only if $G_{k}\models\alpha$ for some $k\in\mathbb{N}$.
@@
Now suppose that $G_{k}\models\alpha$. Define an embedding $f:G_{k}\to G$ such that $f(\bar{\gamma})_{k}=\bar{\gamma}$ and $f(\bar{\gamma})$ is just the identity on all orbits $O_{G,l}$ with $l>k$. Note that since $G$ has orbit independence, we will in fact have that $f(\bar{\gamma})\in G$ and so $f$ is a well defined embedding. By the Łoś-Tarski preservation theorem, existential formulas are preserved under embedding and so $G\models\alpha$. ∎
@@
Let $G$ be a profinite subgroup of $S_{\infty}$ with orbit independence. The existential theory of $G$ is $\Sigma^{0}_{1}$ relative to $\deg(T_{G})$.
@@
There exists a profinite subgroup $G$ of $S_{\infty}$ with orbit independence such that $T_{G}$ is computable and the existential theory of $G$ is $\Sigma^{0}_{1}$ complete.
@@
Let $G$ be any profinite subgroup of $S_{\infty}$ (not necessarily with orbit independence). The existential theory of $G$ is $\Sigma^{0}_{2}$ relative to $\deg(T_{G})$.
@@
There exists a profinite subgroup $G$ of $S_{\infty}$ (without orbit independence) with $T_{G}$ computable such that the existential theory of $G$ is $\Sigma^{0}_{2}$ complete.
@@
Recall that the set $\text{Fin}=\{e\in\mathbb{N}:|W_{e}|<\infty\}$, where $W_{e}$ is the domain of $\Phi_{e}$, the $e$th Turing program, is $\Sigma^{0}_{2}$ complete (see [15], Theorem 4.3.2). Let $\{p_{n}\}_{n\in\mathbb{N}}$ be the sequence of all primes. Given $n\in\mathbb{N}$, define the formula
@@
This, along with Theorem 3.2, ensures that the existential theory of $G$ is $\Sigma^{0}_{2}$ complete.
@@
Let $G$ be a profinite subgroup of $S_{\infty}$ with orbit independence. Let $\alpha$ be any first order sentence in the language of groups. We have $G\models\alpha$ if and only if $G_{k}\models\alpha$ for all sufficiently large $k\in\mathbb{N}$.